A certain radioactive isotope decays at a rate of 0.1% annually. Determine the half-life of this isotope, to the nearest year.
Answer by stanbon(48535)

A certain radioactive isotope decays at a rate of 0.1% annually. Determine the half-life of this isotope, to the nearest year.
------------------------
A(t) = Ao(1-0.001)^t
(1/2)Ao = Ao(0.999)^t
0.999^t = (1/2)
t = [log(1/2) / log(0.999)]
t = 692.80 years
====================
Cheers,
Stan H.